OpenStudy (anonymous):

will medal please help!! Victor will attend college where tuition is $21,000 a year. Books and supplies cost $1800, housing and food are $17,200, transportation is $2700, and other expenses are $1900 a year. He will take out a student loan for $20,000 a year. He has a $500 grant and will earn $2500 a year at a work-study job. Victor's parents will cover the balance. How much do his parents need to contribute? A. $21,600 B. $18,600 C. $4600 D. $1900

OpenStudy (ltrout):

Well lets do some easy math. 21,000 + 1800 + 17,200 + 2,700 + 1,900 = 44,600. He will take out a student loan of 20,000 each year. 44,600 - 20,000 = 24,600. 24,600 - 500 = 24,100. 24,100 - 2,500 = 21,600. theres the answer! Hope it helps!